OpenStudy (anonymous):

is x-[x] defined from all reals to reals ,(where [x] is the greatest integer function ) one-to-one and onto?

OpenStudy (dan815):

@ganeshie8

OpenStudy (shadowlegendx):

dan you fanboy c;

OpenStudy (dan815):

explain this gane

OpenStudy (dan815):

u gota be a fan of gan

ganeshie8 (ganeshie8):

http://www.mathwords.com/h/horizontal_line_test.htm

ganeshie8 (ganeshie8):

if we restrict the domain to (n, n+1] n = natural number then, it will be one-to-one

ganeshie8 (ganeshie8):

n = integer*

Parth (parthkohli):

Recall that\[x - \lfloor x \rfloor = \{x\}\]Where \(\{x\}\) is the fractional part of \(x\). Do you think that no two different real numbers have the same fractional part? Also, do you think that all real numbers can be there in the fractional part of a number?

OpenStudy (anonymous):

@parth i think the ans is no for both the question.am i right??

Parth (parthkohli):

Yes, if you say that the function is \(\mathbb R \to \mathbb R\), then the answer for both should be no.

Parth (parthkohli):

The range of your function is only \([0,1)\) which doesn't cover all real numbers. And \(f(2.5) = f(1.5)\) for example.

OpenStudy (anonymous):

ohk i get it now thanks a lot to all those who helped me...this is really a very good site